OpenStudy (anonymous):

Which is an expression to find the speed in m/s? d = 100 m, t = 20 s

OpenStudy (anonymous):

It's asking you for a way to write those variables so that you get\[\frac{ meters }{ seconds }\]so, since you have\[d=100meters\]\[t=20seconds\]what should you do to get an answer that has the "meter" units on top, divided by the "second" units

OpenStudy (anonymous):

A. 100 - 20 B. 100 + 20 C. 100/20 D. 100 x 20

OpenStudy (anonymous):

What ur saying is D

OpenStudy (ranga):

\[ \Large Speed = \frac{Distance}{Time} = \frac{d}{t} ~m/s \]

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

@Embryo

OpenStudy (anonymous):

It should be C, if you didn't understand my explanation, Ranga provided a nice representation of the answer

OpenStudy (anonymous):

oh I didnt even see ranga's post that really does explain it!